How do you complete a square? Ex: M to the second power - 20m.
Answer by Nate(3500)

m^2 - 20m = 0 Regular Formula
(m^2 - 20m) = 0 Group for Ease
(m - 10)^2 = 100 The m^2 has coefficient as 1 so change to m^1 and take half of (m) ... don't forget to add the number squared! (-10)^2
~>
(m - 10)^2 = 100
m^2 - 20m + 100 = 100